# Eccentricity signatures in LIGO-Virgo-KAGRA’s BNS and NSBH binaries

### v1

@@ III Observations
We present the first eccentricity constraints for six low-mass gravitational-wave events using the newly developed SEOBNRv5EHM waveform model. To ensure robustness, we cross-validate our results using the TEOBResumS-Dali model, which offers a complementary treatment of eccentricity.
Figure 1: Normalized one-dimensional posterior distributions for the orbital eccentricity $e$ at 20 Hz for six low-mass gravitational wave events, using SEOBNRv5EHM and TEOBResumS-Dali depicted in blue and orange for the dominant (2,2) waveform mode (SEOB,TEOB) and with HOMs included depicted in green and red (SEOBHM, TEOBHM). The posteriors have been smoothed using a Gaussian kernel density estimate. The dotted vertical line indicate the bounds of the 90% confidence interval in each panel. For GW200105 the dashed lines represent the 90% credible interval.
III.1 GW170817, GW190425, GW190814, GW200115, and GW230529
@@
For GW170817, we constrain the orbital eccentricity to $e<0.011$ at a reference frequency of 20 Hz. This estimate is consistent with the upper limits reported in Lenon et al. [68] which report a bound of $e<0.024$ at a gravitational wave frequency of 10 Hz, Pal and Nayak [70] with a bound of $e<0.01$ at 20 Hz, and Nitz et al. [71] with a bound of $e<0.02$ at a reference frequency of 30 Hz. For GW190425, we find $e<0.013$ at 20 Hz. To unify the eccentric definitions, we calculate the equivalent eccentricity for a reference frequency of 10 Hz by calculating a minimizing mismatch function [72]. With a mismatch of 0.0036 in comparison to our 20 Hz reference, we obtain the eccentricity to be 0.0464 which is comparable to previous constraints: $e\leq 0.048$ at 10 Hz reported in Lenon et al. [68], and $e\leq 0.007$ at 10 Hz obtained using the SEOBNRE model in Romero-Shaw et al. [73], Cao and Han [74].
Our recovery of mass estimates for GW190814 is consistent with previous analyses [4]. We constrain the eccentricity to be $\sim e<0.063$ for the HOM run using SEOBNRv5EHM. However, we observe that the TEOBResumS-Dali model shows a mild peak near $e\sim 0.03$ in the HOM run, which is not present in the corresponding SEOBNRv5EHM result. While differences between the dominant (2,2) mode and HOM runs can be attributed to the increased sensitivity of higher-order modes—particularly relevant for GW190814, a system known to exhibit strong HOMs, the discrepancy between SEOBNRv5EHM and TEOBResumS-Dali in the HOM runs may reflect waveform systematics. In particular, GW190814’s large mass ratio presents a significant modeling challenge, and waveform inaccuracies at such extreme mass ratios could lead to differences in the inferred eccentricity distributions, as seen in Fig. 1.
For GW200115, we place an upper limit of $e<0.024$ for a reference frequency of 20 Hz, broadly consistent with the constraint of $e<0.02$ also for a reference frequency of 20 Hz reported in de Lluc Planas et al. [16] and the first constraint of $e<0.06$ at 10 Hz reported in Fei and Yang [14]. However, our eccentric posterior distribution is not bimodal, unlike Fei and Yang [14]. For GW230529, we find that our constraint of $e<0.023$ for a reference frequency of 20 Hz using both waveform models is consistent with de Lluc Planas et al. [16] where they place a bound of $e<0.02$ for the same frequency reference.
@@
Table 1: Measured eccentricities and source parameters for low-mass gravitational-wave events analyzed using the TEOBResumS-Dali and SEOBNRv5EHM models, both with and without the inclusion of HOM. The parameters include eccentricity $e$, detector-frame chirp mass $\mathcal{M}$, and mass ratio $q$ $(m_{1}/m_{2}\,,m_{1}>m_{2})$. For eccentricity, we report the 90% upper limit if the 90th percentile is below 0.1, otherwise we report the median with symmetric 90% credible intervals. For GW200105, we always report the median with symmetric 90% credible intervals.
@@ III.2 GW200105
[…]
Our findings are presented in Fig. 3 where all previous results, with the exception of Fei and Yang [14], have been overlaid. All these measurements support the presence of non-zero eccentricity in GW200105 under the assumption of a uniform prior on eccentricity, suggesting the system may not have fully circularized prior to merger. However, some support still exists near zero, leaving open the possibility of a quasicircular origin.
@@
Figure 2: Comparison of eccentricity posteriors for GW200105 using different priors. None of the waveforms include HOM in the plots. The orange curve shows the posterior assuming a uniform prior on eccentricity. The yellow dashed curve corresponds to a posterior assuming a log-uniform prior, with a minimum bound of $e_{\mathrm{min}}=10^{-4}$. The dark blue curve represents the SEOB posterior with a uniform prior, while the dashed dark blue line shows the corresponding posterior under a log-uniform prior with the same lower bound. Vertical dashed lines indicate the median and 90% credible intervals for each corresponding uniform distributions. The 90% credible intervals are depicted in the legend.
@@
While our eccentricity estimates generally align with previous studies, some differences remain. We observe three distinct peaks in our eccentricity posteriors, whereas other works recover two. In particular, de Lluc Planas et al. [16] does not observe the slight peak near $e\sim 0.05$ eccentricity seen in our results and those of Morras et al. [15]. Additionally, Morras et al. [15] does not recover the dual peak structure that we, along with de Lluc Planas et al. [16], observe near $e\sim 0.12$. Both TEOBResumS-Dali and SEOBNRv5EHM agree very well with each other and recover the triple peak distribution. The detector-frame chirp masses are consistent across studies, indicating agreement on the overall signal scale. However, our inferred mass ratio differs significantly from Morras et al. [15], who measure $q=7.63^{+0.32}_{-0.47}$, compared to our SEOBNRv5EHM HOM estimate of $q=3.574^{+0.035}_{-0.030}$. These discrepancies may stem from differences in waveform systematics across multiple waveforms.
@@
Since both our model are non-precessing and assumes aligned spins, it does not capture the full spin dynamics that may be present in the signal, unlike the models employed by Morras et al. [15], though this model is inspiral only and does not capture the full waveform dynamics. These differences highlight the impact of modeling choices and the need for caution when comparing parameter estimates across waveform families, especially when interpreting claims of eccentricity or inferring astrophysical formation scenarios. The development of faithful eccentric and precessing waveform models that capture the full inspiral-merger-ringdown will be essential for robust inferences in future analyses, particularly to help distinguish potential degeneracies between precession and eccentricity.
@@
While previous analyses have provided important constraints on the eccentricity of GW200105 and other NSBH mergers, they have been fundamentally limited by the capabilities of existing waveform models. For example, TaylorF2Ecc includes post-Newtonian (PN) eccentricity effects but is restricted to the inspiral regime and assumes aligned spins, omitting both higher-order PN corrections and precession [75]. PyEFPE, a more recent model, incorporates both spin-induced orbital precession and eccentricity but remains limited to the inspiral phase and does not include merger or ringdown contributions [76]. In contrast, IMRPhenomXP and IMRPhenomXPHM model the full inspiral-merger-ringdown (IMR) evolution and include precession (with the latter adding HOMs), but both are restricted to quasicircular orbits and therefore cannot capture eccentricity [77, 78]. The absence of waveform models that fully capture both eccentricity and the complete IMR evolution—including precession and HOM represents a major limitation of previous studies. The recently introduced IMRPhenomTEHM model attempts to bridge this gap by incorporating eccentric PN dynamics, HOM, and an accurate quasicircular limit, but it remains limited to aligned spins and is still undergoing validation [79]. Notably, IMRPhenomTEHM is comparable to SEOBNRv5EHM for low masses, mass ratios, and low eccentricity, however, exhibits relatively large mismatches when compared with SEOBNRv5EHM in certain regions of parameter space, (larger eccentric values, HOM, large mass ratios) [79], reflecting the fact that, as a newly developed model, further work is required to improve its faithfulness across the full range of astrophysical scenarios.
Figure 3: Corner plot comparing marginalized posterior distributions for GW200105. The parameters shown are chirp mass $\mathcal{M}$, mass ratio $q=m_{1}/m_{2}$ where $m_{1},m_{2}$ are the masses of the primary and secondary object, eccentricity $e$ defined at 20 Hz, aligned spin components $\chi_{1z}$ and $\chi_{2z}$, and mean or relativistic anomaly $\ell$ which is different depending on the model. The distributions are plotted for dominant modes of SEOBNRv5EHM (dominant mode green solid line), SEOBNRv5EHM (HOM blue solid line), TEOBResumS-Dali (dominant mode yellow dash-dot line), and TEOBResumS-Dali (HOM red dash-dot line). The two-dimensional panels display contours representing the 90% credible regions. The one-dimensional histograms along the diagonal show the marginalized posterior distributions for chirp mass, mass ratio, eccentricity, spins and anomaly, with the quoted numbers representing the 90% credible interval. For additional comparison, published results from Morras et al. [15] using pyEFPE (purple dashed line) and de Lluc Planas et al. [16] using IMRPhenomTEHM (pink dotted line) are also included. de Lluc Planas et al. [16] reports the only results with no second island on the eccentricity and chirp mass or mass ratio posteriors, or a mild peak near $e\sim 0.05$. Morras et al. [15] reports the only result with no secondary peak $e\sim 0.12$ in the posterior distribution.
@@ IV Result Verification
Next, we verified the integrity of the waveform generation at higher eccentricities, confirming that the waveforms were free of boundary effects and numerical irregularities. No abnormalities were observed. We also ensured that the waveforms were smoothly tapered to avoid sudden onsets that might mimic eccentric features. Additionally, all likelihood calculations were initialized 1 Hz above the lower frequency cutoff to guard against pseudo-eccentricity effects due to rigid waveform start times.
@@
To asses how much residual eccentricity remains in the less sensitive high-frequency band, we repeated the analysis while computing the likelihood above 30 Hz, and setting the reference and the lower frequency cutoff to 20 Hz. This effectively removes some of the eccentricity information in the early parts of the inspiral, where eccentricity leaves its strongest imprint. The inferred eccentricity remained relatively high at $0.124^{+0.027}_{-0.066}$, which could potentially be due to some underlying noise artifact. To validate this result, we performed a targeted injection study in which we create a fiducial signal of the highest likelihood value from the posteriors of the 30 Hz run. We inferred the source properties of this simulated signal with the same frequency settings as the initial test and recovered an eccentricity of $0.122^{+0.025}_{-0.026}$ consistent with the actual data. This effectively rules out potential noise in the data that could potentially cause large residual eccentricity.
One notable feature in the posterior distribution for GW200105 is the presence of two distinct islands in the eccentricity and chirp mass (or mass ratio) posteriors. This island was also observed by Morras et al. [15]. To investigate this feature, we varied the low-frequency cutoff in the likelihood calculation. As the cutoff was increased from 20 Hz to 30 Hz, the islands in the posterior distribution disappeared. To further assess this effect, we selected a high-likelihood point from the 30 Hz posterior, which initially resided in the valley of the 20 Hz posterior distribution. We recalculated the likelihood at this point and observed an increase as the frequency cutoff was raised. This indicates that the eccentric waveform model effectively rules out posterior samples in the valley, and the appearance of the islands is not due to waveform generation errors, sampling artifacts, or other numerical issues. Furthermore, we also tested different samplers, including Multinest [80], and the islands were also present in this run. Overall, we find that the islands are physical and not an effect of sampling or waveform errors. These islands may be due to the degeneracies between the chirp mass and eccentricities we discuss in Sec. IV.1. Overall, it is still unclear what is causing these two distinct islands.
@@
$\text{SSDR}=\frac{p(d\mid h_{0})}{p(d\mid h_{1})}=\frac{p(e\mid d,h_{1})}{p(e=0\mid h_{1})}\,,$ (2)
where, $p(d\mid h_{0})$ is the marginal likelihood evaluated for a given model $h_{0}$ and $p(d\mid h_{1})$ for the second model, respectively. This ratio provides a direct measure of the Bayes Factor in favor of the eccentricity hypothesis, with values greater than 1 indicating preference for a non-zero eccentricity model.
We computed the Bayes factor in favor of the eccentric model over the non-eccentric model by evaluating the ratio of the posterior to the prior density evaluated at $e=0$. This yielded a Bayes factor of approximately 12.18 for GW200105 using TEOBResumS-Dali and 13.9213.9213.92 using SEOBNRv5EHM. For the HOM we obtain the Bayes factor to be 15.2515.2515.25 and 11.6711.6711.67 for TEOBResumS-Dali and SEOBNRv5EHM, respectively, for a bound ranging from 0 to 0.2 on eccentricity. This indicates roughly 101510-1510 - 15 times more preference for the eccentric model over the non-eccentric one assuming a uniform prior. We also calculate the Bayes factor for a uniform log prior for a dominant mode TEOBResumS-Dali model and we find the value to be 2.352.352.35 in favor of the eccentric model with a bound ranging from $10^{-4}$ to 0.2. The Bayes factors calculated from SDDR are consistent with the evidence reported in Dynesty. While there is still slight support from the uniform log prior run, the evidence is not robust enough to conclusively identify GW200105 as an eccentric binary.
Figure 4: Marginalized posterior distributions and matches for two configurations: one with only eccentricity and anomaly varied (top), and the second with chirp mass and eccentricity varied (bottom). Multiple distinct modes are visible in parameters such as eccentricity, anomaly, and masses, illustrating the complex structure of the likelihood surface. The left plots show the inferred posteriors from parameter estimation, while the right plots show waveform matches for a fixed reference waveform depicted with the red point. The multi modal structure in parameter space reflects underlying features of the waveform evolution. Both TEOBResumS-Dali and SEOBNRv5EHM have this underlying nature in the waveform.

### v2

@@ III Observations
We present the first eccentricity constraints for seven low-mass gravitational-wave events using the newly developed SEOBNRv5EHM waveform model. To ensure robustness, we cross-validate our results using the TEOBResumS-Dali model, which offers a complementary treatment of eccentricity.
Figure 1: Normalized one-dimensional posterior distributions for the orbital eccentricity $e$ at 20 Hz for seven low-mass gravitational wave events, using SEOBNRv5EHM and TEOBResumS-Dali depicted in blue and orange for the dominant (2,2) waveform mode (SEOB,TEOB) and with HOMs included depicted in green and red (SEOBHM, TEOBHM). The posteriors have been smoothed using a Gaussian kernel density estimate. The dotted vertical line indicates the bounds of the 90% confidence interval in each panel. For GW200105, the dashed lines represent the 90% credible interval.
III.1 GW170817, GW190425, GW190814, GW200115, GW230518 and GW230529
@@
For GW170817, we constrain the orbital eccentricity to $e<0.011$ at a reference frequency of 20 Hz. This estimate is consistent with the upper limits reported in Lenon et al. [79] which report a bound of $e<0.024$ at a gravitational wave frequency of 10 Hz, Pal and Nayak [81] with a bound of $e<0.01$ at 20 Hz, and Nitz et al. [82] with a bound of $e<0.02$ at a reference frequency of 30 Hz. For GW190425, we find $e<0.013$ at 20 Hz. To unify the eccentric definitions, we calculate the equivalent eccentricity for a reference frequency of 10 Hz by calculating a minimizing mismatch function [83]. With a mismatch of 0.0036 in comparison to our 20 Hz reference, we obtain the eccentricity to be 0.0464, which is comparable to previous constraints: $e\leq 0.048$ at 10 Hz reported in Lenon et al. [79], and $e\leq 0.007$ at 10 Hz obtained using the SEOBNRE model in Romero-Shaw et al. [84], Cao and Han [85].
Our recovery of mass estimates for GW190814 is consistent with previous analyses [4]. We constrain the eccentricity to be $e<0.063$ for the HOM run using SEOBNRv5EHM. However, we observe that the TEOBResumS-Dali model shows a mild peak near $e\sim 0.03$ in the HOM run, which is not present in the corresponding SEOBNRv5EHM result. Although differences between the dominant (2,2) mode and HOM runs can be attributed to the increased sensitivity of higher-order modes, particularly relevant for GW190814, a system known to exhibit strong HOMs, the discrepancy between SEOBNRv5EHM and TEOBResumS-Dali in the HOM runs may reflect waveform systematics. In particular, GW190814’s large mass ratio presents a significant modeling challenge, and waveform inaccuracies at such extreme mass ratios could lead to differences in the inferred eccentricity distributions, as seen in Fig. 1.
For GW200115, we place an upper limit of $e<0.024$ for a reference frequency of 20 Hz, broadly consistent with the constraint of $e<0.02$ also for a reference frequency of 20 Hz reported in de Lluc Planas et al. [20] and the first constraint of $e<0.06$ at 10 Hz reported in Fei and Yang [18]. However, our eccentric posterior distribution is not bimodal, unlike Fei and Yang [18]. For GW230529, we find that our constraint of $e<0.023$ for a reference frequency of 20 Hz using both waveform models is consistent with de Lluc Planas et al. [20] where they place a bound of $e<0.02$ for the same frequency reference. For GW230518, we find no evidence of eccentricity and place an upper bound of $e<0.04$ for the HOM models.
@@
Table 1: Measured eccentricities and source parameters for low-mass gravitational-wave events analyzed using the TEOBResumS-Dali and SEOBNRv5EHM models, both with and without the inclusion of HOM. The parameters include eccentricity $e$, detector-frame chirp mass $\mathcal{M}$, and mass ratio $q$ $(m_{1}/m_{2}\,,m_{1}>m_{2})$. For eccentricity, we report the 90% upper limit if the 90th percentile is below 0.1; otherwise, we report the median with symmetric 90% credible intervals. For GW200105, we always report the median with symmetric 90% credible intervals.
@@ III.2 GW200105
The first evidence for eccentric features in GW200105 was reported by Fei and Yang [18], who measured an eccentricity of approximately $0.07^{+0.01}_{-0.01}$ at a reference frequency of 20 Hz using a modified version of the IMRPhenomXPHM approximant [85]. Since then, similar studies have estimated the eccentricity parameter for GW200105, further supporting the presence of eccentricity in the signal. For example, Morras et al. [19] reported a median eccentricity of $\sim 0.145^{+0.007}_{-0.097}$ ($90\%$ credible interval) using several waveform models including TaylorF2Ecc, pyEFPE, IMRPhenomXP, and IMRPhenomXPHM [86, 87, 88, 89], evaluated at a reference frequency of 20 Hz. More recently, de Lluc Planas et al. [20] measured the eccentricity as $\leq 0.12^{+0.02}_{-0.03}$ for a reference frequency of 20 Hz using the waveform model IMRPhenomTEHM [90]. In this work, we measure the eccentricity of GW200105 to be $e_{\text{SEOB}}=0.139^{+0.020}_{-0.092}$ using SEOBNRv5EHM and $e_{\text{TEOB}}=0.137^{+0.021}_{-0.092}$ using TEOBResumS-Dali, both evaluated at a reference frequency of 20 Hz. When including HOMs, we obtain $e_{\text{SEOB}}=0.126^{+0.027}_{-0.082}$ and $e_{\text{TEOB}}=0.135^{+0.019}_{-0.088}$. While the inclusion of HOMs slightly reduces the eccentricity estimate in SEOBNRv5EHM, the distribution remains broadly consistent between both waveform models.
Our findings are presented in Fig. 2, where all previous results, except Fei and Yang [18], have been overlaid. All these measurements support the presence of non-zero eccentricity in GW200105 under the assumption of a uniform prior on eccentricity, suggesting the system may not have fully circularized before merger. However, some support still exists near zero, leaving open the possibility of a quasicircular origin.
Figure 2: Corner plot comparing marginalized posterior distributions for GW200105. The parameters shown are chirp mass $\mathcal{M}$, mass ratio $q=m_{1}/m_{2}$ where $m_{1},m_{2}$ are the masses of the primary and secondary object, eccentricity $e$ defined at 20 Hz, aligned spin components $\chi_{1z}$ and $\chi_{2z}$, and mean or relativistic anomaly $\ell$ which is different depending on the model. The distributions are plotted for dominant modes of SEOBNRv5EHM (dominant mode green solid line), SEOBNRv5EHM (HOM blue solid line), TEOBResumS-Dali (dominant mode yellow dash-dot line), and TEOBResumS-Dali (HOM red dash-dot line). The two-dimensional panels display contours representing the 90% credible regions. The one-dimensional histograms along the diagonal show the marginalized posterior distributions for chirp mass, mass ratio, eccentricity, spins, and anomaly, with the quoted numbers representing the 90% credible interval. For additional comparison, published results from Morras et al. [19] using pyEFPE (purple dashed line) and de Lluc Planas et al. [20] using IMRPhenomTEHM (pink dotted line) are also included. de Lluc Planas et al. [20] reports the only results with no second island on the eccentricity and chirp mass or mass ratio posteriors, or a mild peak near $e\sim 0.05$. Morras et al. [19] reports the only result with no secondary peak $e\sim 0.12$ in the posterior distribution.
Figure 3: Comparison of eccentricity posteriors for GW200105 using different priors. None of the waveforms include HOMs in the plots. The orange curve shows the posterior assuming a uniform prior on eccentricity. The yellow dashed curve corresponds to a posterior assuming a log-uniform prior, with a minimum bound of $e_{\mathrm{min}}=10^{-4}$. The dark blue curve represents the posterior on eccentricity with a uniform prior using SEOBNRv5EHM, while the dashed dark blue line shows the corresponding posterior under a log-uniform prior with the same lower bound. Vertical dashed lines indicate the median and 90% credible intervals for each corresponding uniform distribution. The 90% credible intervals are depicted in the legend.
@@
While our eccentricity estimates generally align with previous studies, some differences remain. We observe three distinct peaks in our eccentricity posteriors, whereas other works recover two. In particular, de Lluc Planas et al. [20] does not observe the slight peak near $e\sim 0.05$ eccentricity seen in our results and those of Morras et al. [19]. Additionally, Morras et al. [19] does not recover the dual peak structure that we, along with de Lluc Planas et al. [20], observe near $e\sim 0.12$. Both SEOBNRv5EHM and TEOBResumS-Dali agree very well with each other and recover the triple peak distribution. The detector-frame chirp masses are consistent across studies, indicating agreement on the overall signal scale. However, our inferred mass ratio differs significantly from Morras et al. [19], who measure $q=7.63^{+0.32}_{-0.47}$, compared to our SEOBNRv5EHM HOM estimate of $q=3.574^{+0.035}_{-0.030}$. These discrepancies may stem from differences in waveform systematics across multiple waveforms.
@@
Since both our models are non-precessing and assume aligned spins, it does not capture the full spin dynamics that may be present in the signal, unlike the models employed by Morras et al. [19], though this model is inspiral only and does not capture the full waveform dynamics. These differences highlight the impact of modeling choices and the need for caution when comparing parameter estimates across waveform families, especially when interpreting claims of eccentricity or inferring astrophysical formation scenarios. The development of faithful eccentric and precessing waveform models that capture the full inspiral-merger-ringdown will be essential for robust inferences in future analyses, particularly to help distinguish potential degeneracies between precession and eccentricity.
@@
While previous analyses have provided important constraints on the eccentricity of GW200105 and other NSBH mergers, they have been fundamentally limited by the capabilities of existing waveform models. For example, TaylorF2Ecc includes post-Newtonian (PN) eccentricity effects but is restricted to the inspiral regime and assumes aligned spins, omitting both higher-order PN corrections and precession [86]. PyEFPE, a more recent model, incorporates both spin-induced orbital precession and eccentricity but remains limited to the inspiral phase and does not include merger or ringdown contributions [87]. In contrast, IMRPhenomXP and IMRPhenomXPHM model the full inspiral-merger-ringdown (IMR) evolution and include precession (with the latter adding HOMs), but both are restricted to quasicircular orbits and therefore cannot capture eccentricity [88, 89]. The absence of waveform models that fully capture both eccentricity and the complete IMR evolution—including precession and HOMs represent a major limitation of previous studies. The recently introduced IMRPhenomTEHM model attempts to bridge this gap by incorporating eccentric PN dynamics, HOM, and an accurate quasicircular limit, but it remains limited to aligned spins and is still undergoing validation [90]. Notably, IMRPhenomTEHM is comparable to SEOBNRv5EHM for low masses, mass ratios, and low eccentricity. However, exhibits relatively large mismatches when compared with SEOBNRv5EHM in certain regions of parameter space (larger eccentric values, HOM, large mass ratios) [90], reflecting the fact that, as a newly developed model, further work is required to improve its faithfulness across the full range of astrophysical scenarios.
@@ IV Result Verification
Next, we verified the integrity of waveform generation at higher eccentricities, confirming that the waveforms were free of boundary effects and numerical irregularities. No abnormalities were observed. We also ensured that the waveforms were smoothly tapered to avoid sudden onsets that might mimic eccentric features. Additionally, all likelihood calculations were initialized 1 Hz above the lower frequency cutoff to guard against pseudo-eccentricity effects due to rigid waveform start times.
@@
To assess how much residual eccentricity remains in the less sensitive high-frequency band, we repeated the analysis while computing the likelihood above 30 Hz, and setting the reference and the lower frequency cutoff to 20 Hz. This effectively removes some of the eccentricity information in the early parts of the inspiral, where eccentricity leaves its strongest imprint. The inferred eccentricity remained relatively high at $0.124^{+0.027}_{-0.066}$, which could potentially be due to some underlying noise artifact. To validate this result, we performed a targeted injection study in which we created a fiducial signal with the maximum-likelihood parameters from the posteriors of the 30 Hz run. We inferred the source properties of this simulated signal with the same frequency settings as the initial test and recovered an eccentricity of $0.122^{+0.025}_{-0.026}$ consistent with the actual data. This suggests that such measurements are plausible and expected without invoking non-Gaussian noise.
One notable feature in the posterior distribution for GW200105 is the presence of two distinct islands in the eccentricity and chirp mass (or mass ratio) posteriors. This island was also observed by Morras et al. [19]. To investigate this feature, we varied the low-frequency cutoff in the likelihood calculation. As the cutoff was increased from 20 Hz to 30 Hz, the islands in the posterior distribution merged into one final island. To further assess this effect, we selected a high-likelihood point from the 30 Hz posterior, which initially resided in the valley of the 20 Hz posterior distribution. We recalculated the likelihood at this point and observed an increase as the frequency cutoff was raised. This indicates that the eccentric waveform model effectively rules out posterior samples in the valley, and the appearance of the islands is not due to waveform generation errors, sampling artifacts, or other numerical issues. Furthermore, we also tested different samplers, including Multinest [91], and the islands were also present in this run. Overall, we find that the islands are physical and not an effect of sampling or waveform errors. These islands may be due to the degeneracies between the chirp mass and eccentricities we discuss in Sec. IV.1. Overall, it is still unclear what is causing these two distinct islands.
The accuracy of TEOBResumS-Dali and SEOBNRv5EHM models are not well validated at larger mass ratios, which motivated our original mass ratio prior bound range of $q\ \epsilon\ [1,6]$. To assess the impact of this choice, we performed a follow-up analysis with an expanded prior $q\ \epsilon\ [1,15]$. Overall, we found that our initial measurements were consistent with the new run and our mass ratio agrees with what was measured in previous analysis [20].
@@
$\text{SSDR}=\frac{p(d\mid h_{1})}{p(d\mid h_{0})}=\frac{p(e\mid d,h_{1})}{p(e=0\mid h_{1})}\,$, (2)
where, $p(d\mid h_{0})$ is the marginal likelihood evaluated for a quasi-circular model $h_{0}$ and $p(d\mid h_{1})$ for the eccentric model $h_{1}$ model respectively. This ratio provides a direct measure of the Bayes Factor in favor of the eccentricity hypothesis, with values greater than 1 indicating preference for a non-zero eccentricity model.
We computed the Bayes factor in favor of the eccentric model over the non-eccentric model by evaluating the ratio of the posterior to the prior density evaluated at $e=0$. This yielded a Bayes factor of approximately $6.712$ for GW200105 using SEOBNRv5EHM and 6.454 using TEOBResumS-Dali. For the HOM, we obtain the Bayes factor to be $6.372$ and $7.814$ for SEOBNRv5EHM and TEOBResumS-Dali, respectively, for a bound ranging from 0 to 0.2 on eccentricity. This indicates roughly 676-7 times more preference for the eccentric model over the non-eccentric one, assuming a uniform prior. We also calculate the Bayes factor for a uniform log prior for a dominant mode TEOBResumS-Dali model, and we find the value to be $2.35$ in favor of the eccentric model with a bound ranging from $10^{-4}$ to 0.2. The Bayes factors calculated from SDDR are consistent with the evidence reported in Dynesty. While there is still slight support from the uniform log prior run, the evidence is not robust enough to conclusively identify GW200105 as an eccentric binary.
Figure 4: Marginalized posterior distributions and matches for two configurations: one with only eccentricity and anomaly varied (top), and the second with chirp mass and eccentricity varied (bottom). The model used for the plot is the dominant (2,2) mode SEOBNRv5EHM. The same test was performed for TEOBResumS-Dali and the same distribution is present in both waveforms. Multiple distinct modes are visible in parameters such as eccentricity, anomaly, and masses, illustrating the complex structure of the likelihood surface. The left plots show the inferred posteriors from parameter estimation, while the right plots show waveform matches for a fixed reference waveform depicted with the red point. The multi-modal structure in parameter space reflects underlying features of the waveform evolution.